Mikaela1509:

You mix 1/4 cup of crisp wheat cereal with 1 & 1/2 cups of crisp rice cereal. Find the value of the ratio of crisp wheat cereal to crisp rice cereal. (Hint: Divide the fractions - Copy/Dot/Flip) A: 1/6 B: 1/5 C: 1/2 D: 1/3

TETSXPREME:

do you know how to divide the fraction?

Mikaela1509:

yes and no

TETSXPREME:

well what do you know?

TETSXPREME:

you must divide 2/4 to get the answer for the denominator

snowflake0531:

You mix 1/4 cup of crisp wheat cereal with 1 & 1/2 cups of crisp rice cereal. Find the value of the ratio of crisp wheat cereal to crisp rice cereal. basically you need to find the ratio between 1/4 and 3/2, so simplify \[\frac{\frac{1}{4} }{ \frac{3}{2}} \]